OpenStudy (anonymous):

PLEASE HELPPPPPPPP What is the axis of symmetry for y = x^2 - 8x + 8?

terenzreignz (terenzreignz):

Do you know how to find the vertex of a quadratic function such as this? ^_^

OpenStudy (anonymous):

no. i don't . i do on a graph

terenzreignz (terenzreignz):

Oh, well, in that case, you can use a formula... the vertex for a quadratic equation of the form \[\Large y = \color{red}ax^2+\color{blue}bx+\color{green}c\]is \[\Large \left(-\frac{\color{blue}b}{2\color{red}a}\qquad , \qquad\frac{4\color{red}a\color{green}c-\color{blue}b^2}{4\color{red}a}\right)\]

terenzreignz (terenzreignz):

But the axis of symmetry is just the line \[\huge x = -\frac{\color{blue}b}{2\color{red}a}\]

OpenStudy (anonymous):

so is it -8/x^2

OpenStudy (anonymous):

-b ----- will give you your x value for the turning point and x= bla is your 2a axis of symmetry :D

OpenStudy (anonymous):

so x=4

OpenStudy (anonymous):

what about this one ? What is the axis of symmetry for y = 4x^2 + 12x + 9?

OpenStudy (anonymous):

use the formula -b/2a -> -12/8 = 3/2

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

nvm Your forgot the -